OpenStudy (anonymous):

Suppose you have a mean standarized score of 1500 points with a standard deviation of 175 points. This data is normally distributed. What is the z-score of a 1350 points?

OpenStudy (anonymous):

A. -0.86 B. 0.86 C. 1.54 D. 1.67

OpenStudy (sirm3d):

use the formula \[\large z=\frac{x-\mu}{\sigma}\]

OpenStudy (anonymous):

idk how

OpenStudy (sirm3d):

where x=score that is to be converted to z-score \(\mu\)=mean standard score \(\sigma\)=standard deviation

OpenStudy (anonymous):

so x-1500/175?

OpenStudy (anonymous):

how do i get the x please help

OpenStudy (sirm3d):

x is that score that you should convert to a z-score. check the problem. it says that you convert something to a z-score.

OpenStudy (anonymous):

1.67 =(1750-1500)/150

OpenStudy (anonymous):

THANKS SO MUCH

OpenStudy (anonymous):

Medal? :D (3) z - score = (raw score - mean score) / (standard deviation) z - score = (1750 - 1500) / (175)

OpenStudy (anonymous):

THE ANSWER IS -O.86

OpenStudy (sirm3d):

oops. it says what is the z-score of a 1350 points? @FutureMathProfessor

OpenStudy (sirm3d):

you should try it yourself, @840097 with x=1350.

OpenStudy (anonymous):

I Did i got the answer thanks :)

OpenStudy (sirm3d):

yw